OpenStudy (anonymous):

Write an equation in standard form of the line passing through the points (4,5) and (-4,7) I know standard for is ax+by=c, but I am not sure how to solve this or set it up

TheSmartOne (thesmartone):

Hi! Welcome to OpenStudy! :D

TheSmartOne (thesmartone):

First of all, do you know the slope formula? @KristenSkate

OpenStudy (anonymous):

y=mx+b

TheSmartOne (thesmartone):

That's the slope-intercept form. But do you know the slope formula?

OpenStudy (dan815):

.

OpenStudy (alexandervonhumboldt2):

the slope formula is \[\frac{ y_1-y_2 }{ x_1-x_2 }\] where (x1, y1) and (x2, y2) are your points.

OpenStudy (anonymous):

point slope formula?

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (alexandervonhumboldt2):

the formla used to count slope if called slope formula

TheSmartOne (thesmartone):

No, I was talking about the slope formula. The formula that you use to find the slope when you are given two points of a line. :) Alexander has provided us with the formula :)

OpenStudy (alexandervonhumboldt2):

;)

OpenStudy (alexandervonhumboldt2):

so first lets write the equation in slope-intercept form, and then convert it into standard.

TheSmartOne (thesmartone):

Well, anyways, I found a way to get the answer to your question without having to find the slope formula! :D

OpenStudy (alexandervonhumboldt2):

so we have the slope: (5-7)/(4+4)=-2/8=-1/4

TheSmartOne (thesmartone):

@AlexandervonHumboldt2 Let's guide the user, please :)

OpenStudy (alexandervonhumboldt2):

i'm angry on you tso xD

OpenStudy (anonymous):

okay so I subtract 7-5 on the numerator and then for the denominator I do -4-4?

TheSmartOne (thesmartone):

So if you want to find the linear equation of a line that passes through points \(\sf\Large (x_1,y_1), ~and~(x_2,y_2)\) in the form: \(\sf\Large Ax+By+c = 0\) then you use the formula: \(\sf\large (y_2-y_1)x + (x_2-x_1)y + (x_1y_1 - x_2y_1) = 0\)

TheSmartOne (thesmartone):

Sure, it might seem a little complex at first. But let's break it down, ok? :)

OpenStudy (anonymous):

ok

TheSmartOne (thesmartone):

We are given the points (4,5) and (-4,7), right? :)

OpenStudy (anonymous):

yes so 4 is X1 and -4 is X2 5 is Y1 and 7 is Y2

TheSmartOne (thesmartone):

Correct! :D

TheSmartOne (thesmartone):

So, now we just plug that into the formula :)

TheSmartOne (thesmartone):

Once we do it this way, if you want, we can do it another way too! So that way you don't need to solve it one way! :D

OpenStudy (anonymous):

(5-7)divided(4+4)

TheSmartOne (thesmartone):

Yes, that is how we find the slope :D

OpenStudy (anonymous):

so it = -1/4?

TheSmartOne (thesmartone):

So simplifying that we get: \(\Large\text{Slope } = \frac{ y_2 - y_1 } { x_2 - x_1 } \\\Large= \frac{ 7 - 5}{-4 - 4} \\\Large= \frac{ 2}{-8} \\ \Large= - \frac{1}{4}\)

TheSmartOne (thesmartone):

Correct, that is the slope! :D

OpenStudy (anonymous):

how do I put it in standard form?

TheSmartOne (thesmartone):

Do you know the slope-point form? :)

TheSmartOne (thesmartone):

first we can easily make it in slope-point form, and then convert it to standard form :D

OpenStudy (anonymous):

y-y1=m(x-x1) is that it?

TheSmartOne (thesmartone):

Correct! :D

TheSmartOne (thesmartone):

Ok, so let's choose a point. Which point do you want to use? (4,5) or (-4,7) ?

OpenStudy (anonymous):

(4,5)?

TheSmartOne (thesmartone):

sure!

TheSmartOne (thesmartone):

ok so we need to plug in the point (4,5) and the slope which is -1/4 into it 4 is x1, and 5 is y1 m is the slope which is -1/4

TheSmartOne (thesmartone):

So what do we get when we plug that into y-y1=m(x-x1)

OpenStudy (anonymous):

y-5=-1/4(x-4) is that it?

TheSmartOne (thesmartone):

correct! :D

TheSmartOne (thesmartone):

Now lets make that into standard form :)

TheSmartOne (thesmartone):

\(\sf\Large y - 5 = -\frac{1}{4}(x-4)\) Lets distribute the -1/4 into the parenthesis Remember: a(b-c) = ab - ac

OpenStudy (anonymous):

when distributing -1/4(-4) would it cancel out 4 of the denominator making it 1?

TheSmartOne (thesmartone):

mhmm, actually no :(

TheSmartOne (thesmartone):

\(\sf\Large y - 5 = -\frac{1}{4}(x-4)\) \(\sf\Large y - 5 = (-\frac{1}{4})(x) - (-\frac{1}{4})(4)\) \(\sf\Large y - 5 = -\frac{1}{4}x - (-1)\)

TheSmartOne (thesmartone):

Do you understand? :)

pooja195 (pooja195):

No explain further.

TheSmartOne (thesmartone):

@KristenSkate So, now we need to simplify it further. :)

TheSmartOne (thesmartone):

Lets simplify this: \(\sf\Large -\frac{1}{4} -(-1) \) Do you know what happens when we subtract a negative number?

OpenStudy (anonymous):

i know a negative times a negative is a positive

TheSmartOne (thesmartone):

It would become addition :)

TheSmartOne (thesmartone):

So now we have \(\sf\Large y - 5 = -\frac{1}{4}x +1\) Lets add the -1/4x on both sides :)

TheSmartOne (thesmartone):

And finally, we have to add 5 on both sides :D

OpenStudy (anonymous):

y=-1/4x+6

TheSmartOne (thesmartone):

And now all we need to do is add 1/4x on both sides :D

TheSmartOne (thesmartone):

So now we have \(\sf\Large y + \frac{1}{4}x = -\frac{1}{4}x +\frac{1}{4}x +6\) Simplify it :)

OpenStudy (anonymous):

1/4x+y=6

TheSmartOne (thesmartone):

And that's your answer! :D

TheSmartOne (thesmartone):

That is in the form ax + by = c

TheSmartOne (thesmartone):

Thank you for cooperating @kirstenc73093 :D And Welcome to OpenStudy! :D

OpenStudy (anonymous):

thanks so much I hope I can have you help me again!!

TheSmartOne (thesmartone):

Anytime! :D

TheSmartOne (thesmartone):

Just tag me if you ever need me You can tag me like: @thesmartone

TheSmartOne (thesmartone):

\(\bf\huge~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0

OpenStudy (anonymous):

oh okay. how can I give you a good rating?

TheSmartOne (thesmartone):

Ah, I'm not a Qualified Helper, but I'm just as great as one. So the QH's backed off when they saw me helping. But they're just as great. :P You can rate them by clicking this button: 

TheSmartOne (thesmartone):

That button can be found at the top of your screen :)

TheSmartOne (thesmartone):



OpenStudy (anonymous):

@TheSmartOne you should be a paid tutor! You awesome thanks for being patient with me

TheSmartOne (thesmartone):

Anytime! :D

OpenStudy (jmark):

using slope point form, slope m = 7 - 5/-4-4 = 2/-8 = -1/4 equation, (y - 5) = -1/4(x - 4) or 4(y-5)+x-4 = 4y +x=24.

TheSmartOne (thesmartone):

You're late @jmark 23 days late. And the asker already got the answer too. So, I see no reason for you to have replied to this post.